# Whistle: Data-Efficient Multilingual and Crosslingual Speech Recognition via Weakly Phonetic Supervision

### v1

@@ III Approach
In this section, we describe the three main classes of pre-training and fine-tuning methods for MCL-ASR, i.e., phoneme-based multilingual supervised pre-training (Section III-A), subword-based multilingual supervised pre-training (Section III-B) and multilingual self-supervised pre-training (Section III-C). Figure 1 shows the differences between the three methods. We can see from Figure 1 that similar neural network architectures can be used for the acoustic encoders in all the three methods, which is good for fair comparison.
The input to the acoustic encoder is usually spectral features, obtained from short-time Fourier transform frame by frame, denoted by $x_{1},\cdots,x_{T}\triangleq x_{1:T}$. In DNN-based ASR, the acoustic encoder could be viewed as a non-linear feature extractor, which hopefully can be trained to extract high-level features (or say, representations), more discriminative than the raw spectral features. The output representations from the acoustic encoder are denoted by $h_{1},\cdots,h_{T}\triangleq h_{1:T}$. A popular neural network architecture for the encoder is Conformer [45], which consists of convolution blocks followed by Conformer blocks.
Given acoustic observations $x_{1:T}$, the task of ASR is to find the most likely labels $y_{1},\cdots\,y_{L}\triangleq y_{1:L}$. Different units can be used for labeling $y_{1:L}$, depending on what transcription is used for labeling, phonetic or graphemic, as shown in Table II. Phonemes and subwords are two widely-used labels for MCL-ASR.
In order to promote information sharing between different languages for MCL-ASR, training data from a number of languages, often referred to as seen languages, can be merged to pre-train a multilingual encoder in a supervised fashion, with labels of $y_{1:L}$ given in the form of either phonemes or subwords. Alternatively, the acoustic encoder could be pre-trained over unlabeled data by some self-supervised method, such as wav2vec 2.0 [28], and then be fine-tuned over labeled data in the form of either phonemes or subwords.
III-A Phoneme-based multilingual supervised pre-training
@@
$\begin{split}z_{t}&=W^{T}h_{t}\in\mathbb{R}^{|V|+1}\\ P(\pi_{t}=k|x_{1:T})&=\frac{\exp(z_{t}^{k})}{\sum_{j=1}^{|V|+1}\exp(z_{t}^{j})},k=1,\cdots,|V|+1\end{split}$ (1)
@@
where $W\in\mathbb{R}^{(|V|+1)\times D}$ denotes the weight matrix, and we omit the bias vector in describing the linear layer. The un-normalized outputs $z_{t}$ are often called logits, and $z^{k}_{t}$ denotes the logit corresponding to label $k$.
@@
In phoneme-based multilingual supervised pre-training investigated in this paper, which is called Whistle, we take the union of the phoneme inventories from the seen languages to be the alphabet of labels $V_{\text{multi}}$. The $k$-th row vector from the matrix $W$, denoted by $W(k,:)$, could be viewed as the phoneme embedding for phoneme $k$. The logit for phoneme $k$ at frame $t$ is actually an inner product between the phoneme embedding and the representation vector, $z_{t}^{k}=W(k,:)^{T}h_{t}$.
@@
For recognizing speech from a seen language, the pre-trained encoder together with the phoneme embeddings can be directly used without fine-tuning. Specifically, we build a weighted finite state transducer (WFST) [46], obtained by composing the CTC topology, pronunciation lexicon and word-level n-gram language model, and use WFST-based decoding [47, 21]. While requiring pronunciation lexicons (PROLEX), pre-training with phonetic supervision is more advantageous for information sharing between different languages. In this paper, we relax the requirement of gold-standard human-validated PROLEX and transcripts, by leveraging the LanguageNet G2P models [25]. The LanguageNet G2P models are available for 142 languages. The phonemization procedure in Whistle is detailed in Section IV-B.
@@
For crosslingual speech recognition, denote the phoneme inventory for a new, target language (unseen in pre-training) by $V_{\text{cross}}$. For recognizing speech from the target language, we can initialize a CTC-based model from the pre-trained encoder. The embeddings corresponding to the phonemes in $V_{\text{multi}}\cap V_{\text{cross}}$ are directly copied for initialization. For those phonemes that are not included in the multilingual phoneme alphabet $V_{\text{multi}}$ but appeared in the target language inventory $V_{\text{cross}}$, we randomly initialize their phoneme embeddings. The initialized CTC model can then be fine-tuned over labeled speech from the target language. In this way, the fine-tuned encoder and phoneme embeddings can be used to calculate the logits and the posteriori distribution of $\pi_{t}$ in CTC, and WFST-based decoding can be applied for recognizing speech from the target language.
III-B Subword-based multilingual supervised pre-training
@@
Multilingual supervised pre-training based on subwords is very similar to that based on phonemes, as described in Section III-A, which can still base on the CTC method and use WFST-based decoding with word-level n-gram language model. The major difference is that subword-based multilingual supervised pre-training employs subwords for labeling. Thus, the alphabet of labels $V$ consists of subwords; the lexicon for WFST-based decoing is an orthography lexicon (i.e., words are formed by a sequence of subwords); The row vectors from the matrix $W$ could be viewed as embeddings for subwords. Converting text into subwords is often referred to tokenization, which is still under investigation and needs a balance between granularity and ASR performance [12].
@@
In this paper, we use Byte Pair Encoding (BPE) based subwords, or say, tokens [48]. BPE introduces a word segmentation algorithm, which initializes the token alphabet with the character alphabet and iteratively merges the most frequent pair of tokens. In this way, BPE obtains a compact token vocabulary of variable-length subword units. Notably, the merging of tokens in BPE is based on their frequencies. A straightforward application of BPE may inappropriately favor the merging from high-resource languages; for low-resource languages, tokens may be mostly single characters. Similar to [49], sentences are sampled according to a multinomial distribution with probabilities $\left\{q_{l}\right\}_{l=1...N}$:
$q_{l}=\frac{p_{l}^{\beta}}{\sum_{i=1}^{N}p_{i}^{\beta}}\quad with\quad p_{l}=\frac{n_{l}}{\sum_{i=1}^{N}n_{i}},$ (2)
@@
where $\beta$ controls the sampling of languages with different frequencies. We use $\beta$ = 0.5 in experiments. $N$ is the number of seen languages in the training data, and $n_{l}$ denotes the number of sentences for language $l$. By such data sampling, we can increase the number of tokens associated to low-resource languages and reduce the bias towards high-resource languages.
III-C Multilingual self-supervised pre-training
@@
We pre-train a wav2vec 2.0 model [28] on our multilingual pre-training data (just audio data). The basic architecture of the wav2vec 2.0 model is as follows. A convolutional feature encoder maps raw audio $x_{1:T}$ to latent speech features $z_{1},\dots,z_{T}$, which are then fed to a Transformer to output contextual representations $h_{1},\dots,h_{T}$ [50, 37]. The Transformer architecture is the same as in BERT [51, 37]. During training, a quantization module is employed to discretize the latent features $z_{1},\dots,z_{T}$ to $q_{1},\dots,q_{T}$, which represent the targets in the contrastive learning objective. The quantization module uses a Gumbel softmax to choose entries from the codebooks and the chosen entries are concatenated to be $q_{1},\dots,q_{T}$ [52, 53, 50]. The wav2vec 2.0 model is trained by solving a contrastive task on masked feature encoder outputs. During training, spans of ten time steps with random starting indices are masked. The objective is to predict the true quantized latent $q_{t}$ for masked time-steps within a set of $K=100$ distractors sampled from other masked time steps.
@@
Basically, the pre-trained wav2vec 2.0 model is only an acoustic encoder, consisting of a convolutional feature encoder and a transformer contextual encoder. In order to recognize speech from any language, we need to introduce a linear layer (parameterized by matrix $W$) followed by softmax on top of the encoder output $h_{1},\cdots,h_{T}$, as shown in Eq. (1), and perform fine-tuning over labeled data. The labels could be in the form of either phonemes or subwords.
@@ IV-A Dataset
TABLE I: Multilingual and crosslingual data information, including the language code, the language family, the number of IPA phonemes, the sizes of train, development and test sets in hours.
Code Language Family IPA Hours
Train Dev Test
Multi. en English West Germanic 39 2227.3 27.2 27.0
es Spanish Romance 32 382.3 26.0 26.5
fr French Romance 33 823.4 25.0 25.4
it Italian Romance 30 271.5 24.7 26.0
ky Kyrgyz Turkic 32 32.7 2.1 2.2
nl Dutch West Germanic 39 70.2 13.8 13.9
ru Russian East Slavic 32 149.8 14.6 15.0
sv Swedish North Germanic 33 29.8 5.5 6.2
tr Turkish Turkic 41 61.5 10.1 11.4
tt Tatar Turkic 31 20.8 3.0 5.7
Cross. pl Polish West Slavic 35 129.9 11.4 11.5
id Indonesian Austronesian 35 20.8 3.7 4.1
We conduct experiments on the CommonVoice dataset [54] released at September 2022 (v11.0). CommonVoice is a large multilingual speech corpus, with spoken content taken primarily from Wikipedia articles. We select ten languages for multilingual pre-training experiments: English (en), Spanish (es), French (fr), Italian (it), Kyrgyz (ky), Dutch (nl), Russian (ru), Swedish (sv), Turkish (tr) and Tatar (tt), with a total of 4069.3 hours, which cover rich language families. We refer to this dataset of 10 languages as CV-Lang10. We select Polish (pl) and Indonesian (id) for crosslingual finetuning experiments, which are from two unseen language families. Detailed database descriptions are shown in Table I. We combine all data from the ten languages to form the training, development, and test sets for multilingual pre-training experiments. For each language, we use its transcripts to separately train a word-level n-gram language model for WFST-based decoding.
@@
IV-B Text normalization and phonemization
The FST (Finite State Transducer) based G2P toolkit, Phonetisaurus [26], is utilized to generate labeling of utterances in IPA phonemes from text transcripts. The trained FSTs for use with Phonetisaurus can be obtained from LanguageNet [25]. Examples of phoneme annotations for each language in CV-Lang10 are shown in Table II. By applying Phonetisaurus G2P tool with LanguageNet FSTs, we can also create a PROLEX for each language, which is needed for WFST-based decoding with phoneme-based CTC model. The phonetic transcripts and the PROLEXs for CV-Lang10 will be released in our public repository.
Figure 2: Counts of phonemes (top) and subwords (down) in the CV-Lang10 training set.
@@ IV-C Model training
The CAT toolkit [22] is used for training CTC [4] based ASR models in our experiments. Three sizes of acoustic encoders are used in our experiments, all based on Conformer [45] networks. The small-sized Conformer encoder (S) consists of 14 encoder blocks with dimension 512. We set the self-attention layer to have 4 heads with 36-dimension hidden states, and the feed-forward network (FFN) dimension to 512. The middle-sized Conformer encoder (M) uses 22 blocks, model dimension 640, FFN dimension 640, attention dimension 160, while the large-sized Conformer encoder (L) uses 22 blocks, model dimension 1024, FFN dimension 1024, attention dimension 224. For phoneme-based models, the multilingual alphabet size of phonemes is 73. For subword-based models, the multilingual alphabet size of subwords is 4998. Counting statistics for phonemes and subwords over CV-Lang10 are shown in Figure 2.
@@
TABLE III: Phoneme error rates (PERs) for phoneme-based monolingual models and multilingual pre-trained models on the CV-Lang10 dataset. (S: small, M: middle, L: large)
@@
TABLE IV: Word error rates (WERs) for phoneme-based monolingual models and multilingual pre-trained models on the CV-Lang10 dataset, compared with the subword-based multilingual pre-trained model.
id Model Size (M) en es fr it ky nl ru sv tr tt Avg.

### v2

@@ III Approach
In this section, we describe the three main classes of pretraining and finetuning methods for MCL-ASR, i.e., phoneme-based multilingual supervised pretraining (Section III-A), subword-based multilingual supervised pretraining (Section III-B) and multilingual self-supervised pretraining (Section III-C). Figure 1 shows the differences between the three methods. We can see from Figure 1 that similar neural network architectures can be used for the acoustic encoders in all the three methods, which is good for fair comparison.
The input to the acoustic encoder is usually spectral features, obtained from short-time Fourier transform frame by frame, denoted by $\bm{x}_{1},\cdots,\bm{x}_{T}\triangleq\bm{x}_{1:T}$. In DNN-based ASR, the acoustic encoder could be viewed as a non-linear feature extractor, which hopefully can be trained to extract high-level features (or say, representations), more discriminative than the raw spectral features. The output representations from the acoustic encoder are denoted by $\bm{h}_{1},\cdots,\bm{h}_{T}\triangleq\bm{h}_{1:T}$. A popular neural network architecture for the encoder is Conformer [49], which consists of convolution blocks followed by Conformer blocks.
Given acoustic observations $\bm{x}_{1:T}$, the task of ASR is to find the most likely labels $y_{1},\cdots\,y_{L}\triangleq y_{1:L}$. Different units can be used for labeling $y_{1:L}$, depending on what transcription is used for labeling, phonetic or graphemic, as shown in Table II. Phonemes and subwords are two widely-used labels for MCL-ASR.
In order to promote information sharing between different languages for MCL-ASR, training data from a number of languages, often referred to as seen languages, can be merged to pretrain a multilingual encoder in a supervised fashion, with labels of $y_{1:L}$ given in the form of either phonemes or subwords. Alternatively, the acoustic encoder could be pretrained over unlabeled data by some self-supervised method, such as wav2vec 2.0 [24], and then be finetuned over labeled data in the form of either phonemes or subwords.
III-A Phoneme-based multilingual supervised pretraining
@@
$\begin{split}\bm{z}_{t}&=\bm{W}^{T}\bm{h}_{t}\in\mathbb{R}^{|V|+1}\\ P(\pi_{t}=k|\bm{x}_{1:T})&=\frac{\exp(z_{t}^{k})}{\sum_{j=1}^{|V|+1}\exp(z_{t}^{j})},k=1,\cdots,|V|+1\end{split}$ (1)
@@
where $\bm{W}\in\mathbb{R}^{(|V|+1)\times D}$ denotes the weight matrix, and we omit the bias vector in describing the linear layer. The un-normalized outputs $\bm{z}_{t}$ are often called logits, and $z^{k}_{t}$ denotes the logit corresponding to label $k$.
@@
In phoneme-based multilingual supervised pretraining investigated in this paper, which is called Whistle, we take the union of the phoneme inventories from the seen languages to be the alphabet of labels $V_{\text{multi}}$. The $k$-th row vector from the matrix $\bm{W}$, denoted by $\bm{W}(k,:)$, could be viewed as the phoneme embedding for phoneme $k$. The logit for phoneme $k$ at frame $t$ is actually an inner product between the phoneme embedding and the representation vector, $z_{t}^{k}=\bm{W}(k,:)^{T}\bm{h}_{t}$.
@@
For recognizing speech from a seen language, the pretrained encoder together with the phoneme embeddings can be directly used without finetuning. Specifically, we build a weighted finite state transducer (WFST) [50], obtained by composing the CTC topology, pronunciation lexicon and word-level n-gram language model, and use WFST-based decoding [51, 21]. While requiring pronunciation lexicons (PROLEX), pretraining with phonetic supervision is more advantageous for information sharing between different languages. In this paper, we relax the requirement of gold-standard human-validated PROLEX and transcripts, by leveraging the LanguageNet G2P models [26]. The LanguageNet G2P models are available for 142 languages. The phonemization procedure in Whistle is detailed in Section IV-B.
@@
For crosslingual speech recognition, denote the phoneme inventory for a new, target language (unseen in pretraining) by $V_{\text{cross}}$. For recognizing speech from the target language, we can initialize a CTC-based model from the pretrained encoder. The embeddings corresponding to the phonemes in $V_{\text{multi}}\cap V_{\text{cross}}$ are directly copied for initialization. For those phonemes that are not included in the multilingual phoneme alphabet $V_{\text{multi}}$ but appeared in the target language inventory $V_{\text{cross}}$, we randomly initialize their phoneme embeddings. The initialized CTC model can then be finetuned over labeled speech from the target language. In this way, the finetuned encoder and phoneme embeddings can be used to calculate the logits and the posteriori distribution of $\pi_{t}$ in CTC, and WFST-based decoding can be applied for recognizing speech from the target language.
III-B Subword-based multilingual supervised pretraining
@@
Multilingual supervised pretraining based on subwords is very similar to that based on phonemes, as described in Section III-A, which can still base on the CTC method and use WFST-based decoding with word-level n-gram language model. The major difference is that subword-based multilingual supervised pretraining employs subwords for labeling. Thus, the alphabet of labels $V$ consists of subwords; the lexicon for WFST-based decoing is an orthography lexicon (i.e., words are formed by a sequence of subwords); The row vectors from the matrix $\bm{W}$ could be viewed as embeddings for subwords. In crosslingual finetuning of subword-based pretrained models, we employ the common practice to randomly initialize the parameters in the last linear layer. An ablation study is provided in Section VI-D, which shows that employing the same initialization scheme as in crosslingual finetuning of phoneme-based pretrained models yields worse performance.
@@
Converting text into subwords is often referred to tokenization, which is still under investigation and needs a balance between granularity and ASR performance [12]. In this paper, we use Byte Pair Encoding (BPE) based subwords, or say, tokens [52]. BPE introduces a word segmentation algorithm, which initializes the token alphabet with the character alphabet and iteratively merges the most frequent pair of tokens. In this way, BPE obtains a compact token vocabulary of variable-length subword units. Notably, the merging of tokens in BPE is based on their frequencies. A straightforward application of BPE may inappropriately favor the merging from high-resource languages; for low-resource languages, tokens may be mostly single characters. Similar to [53], sentences are sampled according to a multinomial distribution with probabilities $\left\{q_{l}\right\}_{l=1...K}$:
$q_{l}=\frac{p_{l}^{\beta}}{\sum_{i=1}^{K}p_{i}^{\beta}}\quad with\quad p_{l}=\frac{n_{l}}{\sum_{i=1}^{K}n_{i}},$ (2)
@@
where $\beta$ controls the sampling of languages with different frequencies. We use $\beta$ = 0.5 in experiments. $K$ is the number of seen languages in the training data, and $n_{l}$ denotes the number of sentences for language $l$. By such data sampling, we can increase the number of tokens associated to low-resource languages and reduce the bias towards high-resource languages.
III-C Multilingual self-supervised pretraining
@@
We pretrain a wav2vec 2.0 model [24] on our multilingual pretraining data (just audio data). The basic architecture of the wav2vec 2.0 model is as follows. A convolutional feature encoder maps raw audio $\bm{x}_{1:T}$ to latent speech features $\bm{z}_{1},\dots,\bm{z}_{T}$, which are then fed to a Transformer to output contextual representations $\bm{h}_{1},\dots,\bm{h}_{T}$ [54, 37]. The Transformer architecture is the same as in BERT [55, 37]. During training, a quantization module is employed to discretize the latent features $\bm{z}_{1},\dots,\bm{z}_{T}$ to $\bm{q}_{1},\dots,\bm{q}_{T}$, which represent the targets in the contrastive learning objective. The quantization module uses a Gumbel softmax to choose entries from the codebooks and the chosen entries are concatenated to be $\bm{q}_{1},\dots,\bm{q}_{T}$ [56, 57, 54]. The wav2vec 2.0 model is trained by solving a contrastive task on masked feature encoder outputs. During training, spans of ten time steps with random starting indices are masked. The objective is to predict the true quantized latent $\bm{q}_{t}$ for masked time-steps within a set of $K=100$ distractors sampled from other masked time steps.
@@
Basically, the pretrained wav2vec 2.0 model is only an acoustic encoder, consisting of a convolutional feature encoder and a transformer contextual encoder. In order to recognize speech from any language, we need to introduce a linear layer (parameterized by matrix $\bm{W}$) followed by softmax on top of the encoder output $\bm{h}_{1},\cdots,\bm{h}_{T}$, as shown in Eq. (1), and perform finetuning over labeled data. The labels could be in the form of either phonemes or subwords.
@@ IV-A Dataset
We conduct experiments on the CommonVoice dataset [58] released at September 2022 (v11.0). CommonVoice is a large multilingual speech corpus, with spoken content taken primarily from Wikipedia articles. It is released under a Creative Commons CC0 license, and has often been used in multilingual speech recognition studies [7, 8, 9, 15, 59, 13, 12, 11]. We select ten languages for multilingual pretraining experiments: English (en), Spanish (es), French (fr), Italian (it), Kyrgyz (ky), Dutch (nl), Russian (ru), Swedish (sv), Turkish (tr) and Tatar (tt), with a total of 4069.3 hours, which cover rich language families. We refer to this dataset of 10 languages as CV-Lang10. These ten languages are chosen because they are frequently used to evaluate the performance of multilingual speech recognition systems [7, 15, 59]. We will release the data pre-processing scripts, including text normalization and phonemization, so that people can easily run these scripts themselves to obtain CV-Lang10 from the orignal CommonVoice dataset. Hopefully, in this easy way and based on CC0 license, CV-Lang10 can serve for a common, free experimental setup to facilitate future MCL-ASR research. We select Polish (pl) and Indonesian (id) for crosslingual finetuning experiments, which are from two unseen language families. Detailed database descriptions are shown in Table I. We combine all data from the ten languages to form the training, development, and test sets for multilingual pretraining experiments. For each language, we use its transcripts of training set to separately train a word-level 4-gram language model for WFST-based decoding.
TABLE I: Multilingual and crosslingual data information, including the language code, the language family, the number of IPA phonemes, characters and BPEs for each language.
Code Language Family IPA Char BPE
Multi. en English West Germanic 39 29 3422
es Spanish Romance 32 89 3359
fr French Romance 33 66 3507
it Italian Romance 30 46 3321
ky Kyrgyz Turkic 32 37 784
nl Dutch West Germanic 39 29 2304
ru Russian East Slavic 32 53 974
sv Swedish North Germanic 33 30 2000
tr Turkish Turkic 41 39 1582
tt Tatar Turkic 31 47 773
Cross. pl Polish West Slavic 35 34 500
id Indonesian Austronesian 35 32 500
IV-B Text normalization, phonemization and tokenization
The FST (Finite State Transducer) based G2P toolkit, Phonetisaurus [27], is utilized to generate labeling of utterances in IPA phonemes from text transcripts. The trained FSTs for use with Phonetisaurus can be obtained from LanguageNet [26]. Examples of phoneme annotations for each language in CV-Lang10 are shown in Table II. By applying Phonetisaurus G2P tool with LanguageNet FSTs, we can also create a PROLEX for each language, which is needed for WFST-based decoding with phoneme-based CTC model. The PROLEXs and G2P conversion code for CV-Lang10 will be released in our public repository.
For phoneme-based models, the multilingual alphabet size of phonemes is 73, which can be naturally determined after phonemization. For subword-based systems, we use the BPE tokenizer and empirically determine the multilingual BPE vocabulary size to be 4998 after some pilot experiments. There are two additional special tokens ¡unk¿ and ¡s¿. As explained in Section III-B, when creating the BPE vocabulary for subword-based supervision, training sentences are sampled according to Eq. (2). The number of sentences before and after sampling is shown in Table III for each language. It can be seen that the numbers of sentences for high-resource languages decrease, while those for low-resource languages increase. It can also be seen from Table I that the resulting multilingual subword vocabulary contains a considerable number of subword units in low-resource languages. We can see that this sampling strategy alleviates the problem (to some extent) that the frequencies of subwords are severely biased towards high-resource languages. In summary, the counting statistics for the phoneme and subword units in CV-Lang10 are shown in Figure 2.
@@ IV-B Text normalization, phonemization and tokenization
TABLE III: The number of sentences before and after sampling by Eq. (2) for each language. The sampled sentences are used to train the BPE tokenizer.
en es fr it ky nl ru sv tr tt Total
Before sampling 1,583,721 274,765 607,468 188,038 26,572 61,702 106,294 28,572 62,081 20,352 2,959,565
Sampled 867,689 536,136 361,104 298,887 225,392 172,169 171,133 115,987 112,677 98,391 2,959,565
Figure 2: Counts of phoneme and subword units in the CV-Lang10 training set. Note that this is a log-log plot. The distribution of subwords has a sharp peak around a few top subwords and a severe long tail, which shows a more severe data imbalance than the distribution of phonemes.
@@ IV-C Model training
The CAT toolkit [22] is used for training CTC [4] based ASR models in our experiments. Three sizes of acoustic encoders are used in our experiments, all based on Conformer [49] networks. The small-sized Conformer encoder (S) consists of 14 encoder blocks with dimension 512. We set the self-attention layer to have 4 heads with 36-dimension hidden states, and the feed-forward network (FFN) dimension to 512. The middle-sized Conformer encoder (M) uses 22 blocks, model dimension 640, FFN dimension 640, attention dimension 160, while the large-sized Conformer encoder (L) uses 22 blocks, model dimension 1024, FFN dimension 1024, attention dimension 224. For phoneme-based models, the multilingual alphabet size of phonemes is 73.
@@
TABLE IV: Phoneme error rates (PERs) and Word error rates (WERs) for phoneme-based monolingual models and multilingual pretrained models on the CV-Lang10 dataset, compared with the subword-based multilingual pretrained model. (S: small, M: middle, L: large)
@@
WER